suppose the price elasticity of demand for gasoline is 0.2 in the short run and 0.7 in the long run. if the price of gasoline raise 28 percent, what effect on quantity demanded will this have in the short run? in the long run?
1
Expert's answer
2013-03-13T09:26:51-0400
If you raise the price of gasoline by 28%, demand will change as follows:
in the short term - will decrease by 5.6%, and in the long by 19.6%.


ΔD= - Ep * ΔP
ΔD1= - 0,2 * 28%= - 5,6%
ΔD2= - 0,7 * 28%= - 19,6%